# Flexing RISC-V Instruction Subset Processors (RISPs) to Extreme Edge

### v2

@@ 3.1 Concept of Instruction Hardware Blocks
The RISP generation methodology relies on the RISC-V ISA and its software toolchain. A RISP is automatically generated for a given application or a domain of similar applications, as shown in Figure 2. First, we characterize the application or the domain of applications by compiling them to the target RISC-V ISA (in our case RV32E) to identify the domain-specific instruction subset (see Step 1). The hardware description of each instruction in the instruction subset is called “instruction hardware block” written in RTL (i.e., SystemVerilog), and extracted from a pre-verified full ISA hardware library of all instructions in RISC-V ISA. The pre-verified full ISA hardware library contains “formally verified” instruction hardware blocks, and is, in concept, similar to the standard cell library designed for a process technology used in chip design and implementation, which will be described in Section 3.4.1.
Each instruction in the RISC-V ISA can be thought of as a discrete, fully functional hardware block. Table 2 highlights the block diagram of each instruction type in the RV32I ISA. The semantics of each instruction in the RV32I ISA are converted into its hardware description in RTL, along with a glue logic interfacing with the register file and the memory. The extracted instruction hardware blocks are then automatically combined to build the Modular Execution Unit (ModularEX).
@@ 3.2 Modular Execution Unit
ModularEX is the main execution unit of the generated RISP. It consists of a switch that handles the instruction hardware blocks and their interfaces. The switch is automatically generated and ensures that the outputs of the instruction-under-execution module are triggered. It is a simple case statement in SystemVerilog with N cases, where N is the number of the instructions in the sub-ISA. The instruction hardware blocks are integrated with the switch into ModularEX (fully combinational design).
In Step 3 - Figure 2, ModularEX is stitched to the memory interfaces and the register file (RF) to generate an unoptimised RISP, which will be an input to the hardware synthesis tool to optimize it by maximizing sharing inside ModularEX. The output of this stage is the final RISP tailored to the application or the domain under consideration, which is fully programmable within its domain and does not require any prior design of fundamental blocks used in traditional processors, e.g., functional units like ALUs.
3.3 RISP Construction
Figure 3 shows the construction of a RISP by stitching ModularEX with the rest of the fixed units such as the fetch unit and the RF as well as memory interfaces. By construction, a RISP has a single-cycle processor microarchitecture where an instruction is fetched from the instruction memory and sent to the ModularEX unit, and then, the switch selects the actual instruction hardware block of the fetched instruction executing it at the same cycle. In other words, the switch performs as a partial decoder that selects which block will be enabled at every clock cycle. The full decoding of the instruction is performed inside each instruction block.
Figure 3: A single-cycle RISP microarchitecture created by stitching together ModularEX, the fixed units (i.e., Fetch and RF) and Memory Interfaces.
The key benefit of the RISP generation methodology is that it does not require any expertise in microarchitecture, RTL design or gate netlists. The only input of the framework is the domain-specific instruction subset identified by the application or the domain. At the design level, the user does not need to optimize the microarchitecture manually. Instead, we assign all the optimization decisions to the hardware synthesis tool. The decision of allowing the synthesis tool to remove the redundancies rather than a microarchitect designing a shared unit like an ALU is a conscious one because contemporary commercial/non-commercial hardware synthesis tools are very efficient to discover redundancies and eliminate them. For instance, the synthesis tool will optimize the gate netlists by maximizing the resource sharing if multiple instruction hardware blocks have common operations among them (e.g., design flattening).
@@ 3.4 Verification
Figure 4: The development of the pre-verified full ISA hardware library.
@@ 3.4.1 Development of the Pre-verified Full ISA Hardware Library
Generating a RISP using the novel concept of a pre-verified ISA hardware library significantly reduces the verification complexity of RISP. The full ISA hardware library has been extensively verified in multiple ways including formal verification. It is important to note that the development of the pre-verified full ISA hardware library is a one-time process, and the library can be reused to generate various RISPs from the same ISA based on the requirements of the target domain.
The Full ISA hardware library is a collection of instruction hardware blocks written in SystemVerilog. The library supports the RISC-V RV32I/E instruction set and is fully extendable to support other groups of RISC-V instructions or even custom instructions. Each instruction hardware block is formally verified against its specification in the ISA. This is a significant step towards instruction modularity in the RISP design and its simplified verification. One of the key advantages of this approach is that verification of instruction hardware blocks is integrated into the RISP development process. Every instruction hardware block is formally pre-verified before it is used for the generation of a RISP.
@@
To further extend our verification framework of the library, each testbench has been evaluated with MCY tool from YosysHQ suite [44]. MCY is one of the most popular tools for verification primarily because it offers the ability to generate several mutations of the target RTL design (in our case the implemented instruction blocks). The tool automatically selects only the mutated designs that can cause an important change in the output of the design by filtering them using formal verification techniques. The final step is to automatically check that the testbench can detect and fail for the mutations that can cause an important change in the output of the target design. Under this environment, we ensure that our testbenches are capable of verifying the target instruction hardware blocks.
In terms of formal verification, the full ISA library is formally verified using SymbiYosys [45] (sby), a frontend for Yosys for Formal Verification. For every instruction hardware block, a set of SVA have been developed to ensure that the instruction semantics are satisfied. Also, additional assertions are used to validate the behavior of the interfaces. This approach simplifies the verification process of a RISP, as each instruction hardware block has already been independently and formally verified. Moreover, the process benefits from the well-defined semantics of each instruction in the ISA.
Instruction hardware block representations of all instructions in the ISA are functionally and formally verified, they are stored in the pre-verified full ISA hardware library. When constructing a RISP, instruction hardware blocks from the library will be stitched together to build the ModularEX that does not require any further verification. Therefore, the pre-verified full ISA hardware library concept is similar to the concept of standard cell libraries in VLSI design where the standard cell library for a foundry process is created once and re-used in many chip designs. Each cell is a fully verified building block in a finer granularity at the gate level.
@@ 3.4.2 Integration-level Verification
The pre-verified ModularEX is integrated into the pre-verified fixed units such as the fetch and register file to construct the RISP. Various verification processes have been developed to ensure the correctness of the integrated RISP. They include custom domain checks at various points, such as checks at instruction and register level to verify that each instruction is fetched and executed correctly (custom assembly programs with expected behavior). Additionally, RISC-V compatibility was tested with RISCOF, an architectural test framework [37]. RISCOF automatically chooses a set of test cases for the RV32E ISA, the RISP executes these test cases, and writes a signature to a file. Then, this signature is compared with a signature of a reference model that has been obtained from a RISC-V simulator, such as Spike [38].
Figure 5: Results of instruction profiling and characterization of the three extreme edge applications, and the benchmarks in Embench using different compiler optimization flags. The black dot symbols represent the average number of distinct instructions from the RV32E ISA (shown on the right y-axis) with minimum/maximum values (lines) across different optimization flags.
@@ 4 Experimental Evaluation
One of the key challenges in evaluating the RISP generation methodology is the lack of a representative benchmark suite for extreme edge. We select two extreme edge applications previously demonstrated in the FlexIC technology [26] [29], and also an ML kernel XGBoost (eXtreme Gradient Boosting) [9] as one of the key ML algorithms used in many classification tasks as an extreme edge kernel.
Extreme edge application 1: Armpit malodour classification or armpit [26] is a real-world application in the domain of deodorant industry that advocates the use of flexible organic sensors to detect volatile organic components in odour, and a sensor readout interface and machine learning hardware implemented as a custom FlexIC to score malodour in a human armpit. It is a representative example of a natively flexible smart system [11] consisting of all flexible components from sensors to processing. Two decision trees (one for each gender) are implemented in hardware to classify armpit malodour. The decision trees in the armpit application are used as the benchmark in this paper. The C code of the decision trees was compiled to RV32E to generate a RISP for the application.
Extreme edge application 2: Atrial fibrillation (AF) detection or af_detect [29] is another real-world application in the domain of ECG that uses flexible electronics to detect AF events in a wearable ECG patch. An AF prediction algorithm called “Approximate Pair Presence Tracking” or APPT was implemented as a custom FlexIC. APPT has three stages: i) detection of R peaks in the ECG signal; ii) calculation of RR intervals (the interval between two R peaks) and delta of RR intervals; iii) a Bloom filter-based binary predictor (AF versus non-AF) using a RR and delta RR interval map. The APPT algorithm in af_detect is used as the benchmark in this paper, which is compiled for RV32E to generate a RISP for it.
Extreme edge kernel: XGBoost [9] is known for its performance in classification and regression problems. It is a more accurate and scalable implementation of gradient boosting. For datasets and problems with simpler requirements, decision trees can be mapped on edge devices in the form of C/C++ or on custom hardware implementation processing units to perform ML classification problems, e.g., armpit. We use XGBoost to extract the decision trees from datasets and then implement the generated decision trees in C/C++ to evaluate their RISP requirements. It can be considered more like a kernel in extreme edge applications rather than a full application itself. For evaluation purposes, we use the Pima Indian Diabetes Database [40], a binary-valued data set that predicts whether a patient shows signs of diabetes according to the criteria of the World Health Organization. The dataset consists of 768 instances and 8 attributes.
@@
(a) Maximum clock frequencies in kHz.
(b) NAND2-equivelant gate count.
(c) Total power (static + dynamic).
Figure 6: Performance, power, area (PPA) metrics of custom RISPs (i.e., a RISP per application) compared to two baselines: a RISP supporting the full RV32E instruction set shown as RISP-RV32E, and Serv collected after hardware synthesis.
4.1 Instruction Profile and Analysis of Applications
@@
Before generating a RISP for each application, we need to characterize the applications to identify the distinct instructions in the RV32E instruction set required for the implementation of each RISP. The next step is to explore how the number of distinct instructions is related to the codesize for different compiler optimization flags.
@@
Figure 5 presents the profiling results of the applications from Embench and the three extreme edge applications. The applications have been compiled for RV32E instruction set as baremetal with different compiler optimization flags (--O0, --O1, --O2, --O3, --Oz). Through this analysis, we identify the optimal code size with respect to the distinct instructions for each application for the generation of RISP. The number of distinct instructions ranges from 6 to 32 instructions, and the geometric mean (GEOMEAN) is 18 instructions for different optimization flags across all the applications. Since the RV32E ISA has around 40 instructions, applications use only 31-84% of the full ISA. For the extreme edge applications, these numbers are 24% (XGBoost), 30% (armpit) and 60% (af_detect), respectively. In terms of total number of static instructions across the different compiler optimization results, the GEOMEAN across all the applications are 1228, 677, 687, 897 and 601 for --O0, --O1, --O2, --O3, and --Oz, respectively. For the extreme edge applications, we observe a substantial reduction in static instructions, as we progress from --O0 to --0z. For the extreme edge applications, using the --O2 optimization level flag, we observe 21%, 45.8% and 52% less static instructions than -O0 for armpit, XGBoost and af_detect, respectively, thus, the difference in the number of static instructions between --O2 and --Oz is negligible.
@@
These results indicate that there is an opportunity to improve power/area of a 32-bit RISC-V processor by eliminating unnecessary instructions and designing a customized RISP for a target application or specific domain. Given that the -O2 compiler optimization flag results in minimal discrepancies in average code size with respect to distinct instructions, we select it as the basis for subsequent RISP design for power, area and frequency analysis.
4.2 Power, Area and Frequency Analysis
The RTL of each RISP is synthesized using Pragmatic’s 0.6µm Indium-Gallium-Zinc-Oxide (IGZO)-based FlexIC process using a commercial EDA tool. Each RISP is synthesized without the register file to better understand the effects of the instruction subsets in hardware.

### v3

@@ 3.1 Concept of Instruction Hardware Blocks
The RISSP generation methodology relies on the RISC-V ISA and its software toolchain. A RISSP is automatically generated for a given application or a domain of similar applications, as shown in Figure 2. It starts with the development of the pre-verified full ISA hardware library (see Step 0). This is a one-time effort for a given ISA. The pre-verified full ISA hardware library contains “formally verified” instruction hardware blocks, and is, in concept, similar to the standard cell library designed for a process technology used in chip design and implementation, which will be described in Section 3.4.1. Next, we characterize the application or the domain of applications by compiling them to the target RISC-V ISA (in our case RV32E) to identify the domain-specific instruction subset (see Step 1). The hardware description of each instruction in the instruction subset is called “instruction hardware block” written in RTL (i.e., SystemVerilog), and extracted from a pre-verified full ISA hardware library of all instructions in RISC-V ISA.
Each instruction in the RISC-V ISA can be thought of as a discrete, fully functional hardware block. Table 2 highlights the block diagram of each instruction type in the ISA. The semantics of each instruction is converted into its hardware description in RTL, along with a glue logic interfacing with the Register File (RF) and the memory. The extracted instruction hardware blocks are then automatically combined to build the Modular Execution Unit (ModularEX).
@@ 3.2 Modular Execution Unit
ModularEX is the main execution unit of the generated RISSP. It consists of a switch that handles the instruction hardware blocks and their interfaces. The switch is automatically generated and ensures that the outputs of the instruction-under-execution module are triggered. It is a simple case statement in SystemVerilog with N cases, where N is the number of the instructions in the sub-ISA. The instruction hardware blocks are integrated with the switch into ModularEX (fully combinational design).
In Step 3 - Figure 2, ModularEX is stitched to the memory interfaces and the RF to generate an unoptimised RISSP, which is an input to the hardware synthesis tool that optimizes it by maximizing sharing inside ModularEX. The output is the final RISSP tailored to the application or the domain under consideration, which is fully programmable within the boundaries of its instruction subset.
3.3 RISSP Construction
Figure 3 shows the construction of a RISSP by stitching ModularEX with the rest of the fixed units such as the fetch unit and the RF as well as memory interfaces. By construction, a RISSP has a single-cycle processor microarchitecture where an instruction is fetched from the instruction memory using a 32-bit Program Counter (PC), and sent to the ModularEX unit. Then, the switch selects the actual instruction hardware block of the fetched instruction executing it at the same cycle. In other words, the switch performs as a partial decoder that selects which block will be enabled at every clock cycle. The full decoding of the instruction is performed inside each instruction block.
Figure 3: A single-cycle RISSP microarchitecture created by stitching together ModularEX, the fixed units (i.e., Fetch and RF) and Memory Interfaces
The key benefit of the RISSP generation methodology is that it does not require any expertise in microarchitecture, RTL design or gate netlists. The only input of the framework is the domain-specific instruction subset identified by the application or the domain. All design optimization decisions are taken by the synthesis tool rather than optimizing the design manually. This is a conscious decision because contemporary commercial/non-commercial hardware synthesis tools are very efficient in discovering redundancies and eliminating them. For instance, the synthesis tool will optimize the gate netlists by maximizing the resource sharing if multiple instruction hardware blocks have common operations among them (e.g., design flattening). Any logic optimization performed during synthesis will be formally verified using equivalence checking by the synthesis tool. This approach may not lead to "minimal area/power" but minimizes the design and verification time.
@@ 3.4.1 Development of the Pre-verified Full ISA Hardware Library
Generating a RISSP using the novel concept of a pre-verified ISA hardware library significantly reduces the verification complexity of RISSP. The full ISA hardware library has been extensively verified in multiple ways including formal verification. The development of the pre-verified ISA hardware library is a one-time verification effort that becomes the non-recurring-engineering (NRE) cost of RISSP. The NRE is amortized by the number of RISSPs shipped during the lifetime of this library. In contrast, any new design in conventional CPU/ASIP methodologies requires design and verification, i.e., verification effort becomes a recurring cost.
Figure 4: The development of the pre-verified full ISA hardware library
The full ISA hardware library is a collection of instruction hardware blocks written in SystemVerilog. The library supports the RISC-V RV32I/E instruction set and is fully extendable to support other groups of RISC-V instructions or even custom instructions. Each instruction hardware block is formally verified against its specification in the ISA. This is a significant step towards instruction modularity in the RISSP design and its simplified verification. One of the key advantages of this approach is that verification of instruction hardware blocks is integrated into the RISSP development process. Every instruction hardware block is formally pre-verified before it is used for the generation of a RISSP.
@@
To further extend our verification framework of the library, each testbench is evaluated with MCY tool from YosysHQ suite [17]. MCY is one of the most popular tools for verification primarily because it offers the ability to generate several mutations of the target RTL design (in our case the implemented instruction blocks). The tool automatically selects only the mutated designs that can cause an important change in the output of the design by filtering them using formal verification techniques. The final step is to automatically check that the testbench can detect and fail for the mutations that can cause an important change in the output of the target design. Under this environment, we ensure that our testbenches are capable of verifying the target instruction hardware blocks.
In terms of formal verification, the full ISA library is formally verified using SymbiYosys [14] (sby), a frontend for Yosys for Formal Verification. For every instruction hardware block, a set of SVA are developed to ensure that the instruction semantics are satisfied. Also, additional assertions are used to validate the behavior of the interfaces. This approach simplifies the verification process of a RISSP, as each instruction hardware block is already independently and formally verified. Moreover, the process benefits from the well-defined semantics of each instruction in the ISA.
Instruction hardware block representations of all instructions in the ISA are functionally and formally verified and are stored in the pre-verified full ISA hardware library. When constructing a RISSP, instruction hardware blocks from the library will be stitched together to build the ModularEX that does not require any further verification. Therefore, the pre-verified full ISA hardware library concept is similar to the concept of standard cell libraries in VLSI design where the standard cell library for a foundry process is created once and re-used in many chip designs. Each cell is a fully verified building block in a finer granularity.
@@ 3.4.2 Integration-level Verification
The pre-verified ModularEX is integrated into the pre-verified fixed units, such as fetch and register file to construct the RISSP. Various verification processes and unit-tests have been developed to ensure the correctness of the integrated RISSP. They include custom domain checks at various points, such as checks at instruction and register level to verify that each instruction is fetched and executed correctly. Additionally, RISC-V compatibility is tested with RISCOF, an architectural test framework [16]. RISCOF automatically chooses a set of test cases for the RV32E ISA, the RISSP executes these test cases, and writes a signature to a file. Then, this signature is compared with a signature of a reference model that has been obtained from a RISC-V simulator, such as Spike [37].
In terms of formal verification at processor-level, RISSP has been verified up to a specific depth with riscv-formal [13]. This framework is an open-source standard for the formal verification of RISC-V processors and is based on the RISC-V Formal Interface (RVFI), a standard interface for processors to communicate and verification tools. Specifically, RISSP implements the RVFI interface that mainly tracks valid instructions, register file data and addresses, PC changes and memory accesses. With riscv-formal, we perform various and extensive checks at instruction level for correct execution, checks to ensure that the registers maintain consistency, and PC checks that verify the correct PC values before and after the execution of an instruction.
Figure 5: Results of instruction profiling and characterization of the three extreme edge applications, and the benchmarks in Embench using different compiler optimization flags. The top diagram presents the codesize in KBytes and the bottom diagram illustrates the number of distinct instructions from the RV32E ISA across different optimization flags.
@@ 4 Experimental Evaluation
One of the key challenges in evaluating the RISSP generation methodology is the lack of a representative benchmark suite for extreme edge. We select two extreme edge applications previously demonstrated in the FlexIC technology [29] [32], and also an ML kernel XGBoost (eXtreme Gradient Boosting) [9] as one of the key ML algorithms used in many classification tasks as an extreme edge kernel.
Extreme edge application 1: Armpit malodour classification or armpit [29] is a real-world application in the domain of deodorant industry that advocates the use of flexible organic sensors to detect volatile organic components in odour with a sensor readout interface and machine learning hardware implemented as a custom FlexIC to score malodour in a human armpit. It is a representative example of a natively flexible smart system [11] consisting of all flexible components from sensors to processing. Two decision trees (one for each gender) are implemented in hardware to classify armpit malodour. The decision trees in the armpit application are used as the benchmark in this paper. The C code of the decision trees was compiled to RV32E to generate a RISSP for the application.
Extreme edge application 2: Atrial fibrillation (AF) detection or af_detect [32] is another real-world application in the domain of ECG that uses flexible electronics to detect AF events in a wearable ECG patch. An AF prediction algorithm called “Approximate Pair Presence Tracking” or APPT was implemented as a custom FlexIC. APPT has three stages: i) detection of R peaks in the ECG signal; ii) calculation of RR intervals (the interval between two R peaks) and delta of RR intervals; iii) a Bloom filter-based binary predictor (AF versus non-AF) using a RR and delta RR interval map. The APPT algorithm in af_detect is used as the benchmark in this paper, which is compiled for RV32E to generate a RISSP.
Extreme edge kernel: XGBoost [9] is known for its performance in classification and regression problems. It is a more accurate and scalable implementation of gradient boosting. For datasets and problems with simpler requirements, decision trees can be mapped on edge devices in the form of C/C++ or on custom hardware implementation processing units to perform ML classification problems, e.g., armpit. We use XGBoost to extract a decision tree from a dataset and then implement the generated decision tree in C/C++ to evaluate its RISSP requirements. It can be considered as a kernel in extreme edge applications rather than a full application itself. For evaluation purposes, we use the pima indian diabetes database [39], a binary-valued data set that predicts whether a patient shows signs of diabetes. The dataset consists of 768 instances and 8 attributes.
@@
Table 3: List of distinct instructions per application when compiled with -O2 option
Application List of Distinct Instructions
aha-mont64 [add, addi, and, andi, beq, bge, bgeu, bltu, bne, jal, jalr, lui, lw, or, slli, sltiu, sltu, srai, srli, sub, sw, xor, xori]
crc32 [add, addi, andi, bge, bne, jal, jalr, lui, lw, slli, sltiu, srli, sub, sw, xor, xori]
cubic [addi, and, andi, beq, bge, blt, bne, jal, jalr, lui, lw, slti, sltiu, sw, xor]
edn [add, addi, andi, beq, bge, bne, jal, jalr, lh, lhu, lui, lw, sh, slli, sltiu, sra, srai, srli, sub, sw]
huffbench [add, addi, and, andi, beq, bge, bgeu, blt, bltu, bne, jal, jalr, lbu, lui, lw, or, ori, sb, sll, slli, sltiu, srai, srli, sub, sw]
matmult-int [add, addi, bge, bne, jal, jalr, lui, lw, slli, sltiu, sw]
md5sum [add, addi, and, andi, beq, bge, bgeu, blt, bltu, bne, jal, jalr, lui, lw, or, sb, sll, slli, sltiu, srl, srli, sub, sw, xor, xori]
minver [add, addi, and, beq, bge, bne, jal, jalr, lui, lw, slli, slti, sltiu, sub, sw, xor]
nbody [add, addi, and, andi, beq, bge, bne, jal, jalr, lui, lw, slli, slti, sltiu, srli, sw]
nettle-aes [add, addi, and, andi, beq, bge, bgeu, bltu, bne, jal, jalr, lbu, lui, lw, or, sb, slli, sltiu, srli, sub, sw, xor]
nettle-sha256 [add, addi, and, andi, beq, bge, bgeu, bltu, bne, jal, jalr, lbu, lhu, lui, lw, or, sb, slli, sltiu, sltu, srli, sub, sw, xor]
nsichneu [add, addi, beq, bge, blt, bne, jal, jalr, lui, lw, slli, sltiu, sub, sw]
picojpeg [add, addi, and, andi, beq, bge, bgeu, blt, bltu, bne, jal, jalr, lb, lbu, lh, lhu, lui, lw, or, sb, sh, sll, slli, sltiu, sltu, sra, srai, srli, sub, sw, xori]
primecount [add, addi, beq, bge, blt, bne, jal, jalr, lui, lw, slli, sltiu, sw]
qrduino [add, addi, and, andi, beq, bge, bgeu, blt, bltu, bne, jal, jalr, lbu, lhu, lui, lw, or, ori, sb, sh, slli, sltiu, sltu, sra, srai, srl, srli, sub, sw, xor, xori]
sglib-combined [add, addi, andi, beq, bge, bgeu, blt, bltu, bne, jal, jalr, lbu, lh, lui, lw, sb, sh, slli, sltiu, sltu, srai, sub, sw, xori]
slre [add, addi, and, andi, beq, bge, bgeu, blt, bltu, bne, jal, jalr, lbu, lui, lw, or, slli, slt, sltiu, sltu, srai, sub, sw, xori]
st [add, addi, and, bge, blt, bne, jal, jalr, lui, lw, slli, slti, sltiu, sw]
statemate [addi, beq, bge, blt, bne, jal, jalr, lbu, lui, lw, or, sb, sh, sltiu, sub, sw]
tarfind [add, addi, andi, beq, bge, bgeu, bltu, bne, jal, jalr, lbu, lui, lw, sb, slli, sltiu, srli, sub, sw]
ud [add, addi, beq, bge, blt, bne, jal, jalr, lui, lw, or, slli, sltiu, sub, sw]
wikisort [add, addi, andi, beq, bge, blt, bne, jal, jalr, lui, lw, or, slli, slt, sltiu, sltu, srai, srli, sub, sw]
armpit [add, addi, andi, beq, bge, blt, bne, jal, jalr, lbu, lui, lw, slli, sltiu, sw]
xgboost [addi, andi, bge, blt, jal, jalr, lui, lw, srli, sw, xor, xori]
af_detect [add, addi, andi, beq, bge, bgeu, blt, bltu, bne, jal, jalr, lbu, lui, lw, sb, sh, slli, sltiu, srai, srli, sub, sw, xor]
4.1 Instruction Subset Profile of Applications
@@
Before generating a RISSP for each application, we need to characterize each application to identify the distinct instructions from the RV32E instruction set required to implement each RISSP. The next step is to explore how the number of distinct instructions is related to the codesize for different compiler optimization flags.
@@
Figure 5 presents the profiling results of the applications from Embench and the three extreme edge applications. The applications have been compiled for RV32E instruction set as baremetal (without support of stdlib, libc, libgcc and startfiles) with different compiler optimization flags (--O0, --O1, --O2, --O3, --Oz) using the standard riscv32-gnu-toolchain (version 13.2.0). Through this analysis, we identify the optimal code size with respect to the distinct instructions for each application for the generation of RISSP. The number of distinct instructions ranges from 9 to 32 instructions, and the average is 19 instructions for different optimization flags across all applications. Since the RV32E ISA has around 40 instructions, applications use only 24-86% of the full ISA. For the extreme edge applications, these numbers are 42% (armpit) 31% (xgboost), and 61% (af_detect). In terms of total number of static instructions across the different compiler optimization results, the average across all the applications are 2027, 1149, 1207, 1586 and 1018 for --O0, --O1, --O2, --O3, and --Oz, respectively. For the extreme edge applications, we observe a substantial reduction in static instructions, as we progress from --O0 to --0z. Specifically, using the --O2 optimization level flag, we observe 75%, 74% and 69% less static instructions (reduction in codesize) than -O0 for armpit, xgboost and af_detect, respectively, thus, the difference in the number of distinct instructions between --O2 and --Oz is negligible.
@@
These results indicate that there is an opportunity to improve power/area of a 32-bit RISC-V processor by eliminating unnecessary instructions and designing a customized RISSP for a target application. Given that the -O2 compiler flag results in minimal discrepancies in average code size with respect to distinct instructions, we select it as the basis for subsequent RISSP design for power, area and frequency analysis. Table 3 lists the number of distinct instructions per application when compiled with -O2 option.
4.2 FlexIC Synthesis Results